OpenStudy (anonymous):

Which is the 25th term of the arithmetic sequence where a1 = 7 and a9 = 71 ?

OpenStudy (anonymous):

the general formula is: an=a1+(n-1)d from this equation you can determine d then determine a25

OpenStudy (anonymous):

can you give me the answer? I am checking to see if i got it right. my answer is 151.

OpenStudy (anonymous):

@ghass1978

OpenStudy (anonymous):

There are 8 equal steps between 7 and 71. What is the step size?

OpenStudy (anonymous):

What are you talking about? I just need the answer...to check. Am I right or wrong?

OpenStudy (anonymous):

you're wrong d=8 so a25=7+24.8=199

OpenStudy (anonymous):

okay thank you.